# A Unique Bosonic Symmetry in a 4D Field-Theoretic System

### v1

@@ 4 Bosonic Symmetry Transformations: Uniqueness Property and the Full Set of CF-Type Restrictions
Thus, ultimately, we have been able to establish the uniqueness of the bosonic symmetry transformation operator at the level of symmetry transformations on (i) the individual fields [cf. Eq. (27)], and (ii) the coupled (but equivalent) Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ [cf. Eqs. (31),(32)]. We would like add here that this key observation is true only on the space of quantum fields where all the four CF-type restrictions are satisfied. It is interesting to end this section with the final remark that the unique (i.e. $s_{\omega}=-\,s_{\bar{\omega}}$) bosonic symmetry transformation operator commutes (i.e. $s_{\omega},\;s_{(a)b}]=0,\;[s_{\omega},\;s_{(a)d}]=0$) with all the four nilpotent symmetry transformation operators (i.e. $s_{(a)b},\;s_{(a)d}$) of our theory.
5 Conclusions
@@
In our present endeavor, we have very briefly sketched the existence of the infinitesimal, continuous and off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,s_{(a)d}^{2}=0$) versions of the (anti-)BRST and (anti-)co-BRST symmetry transformation operators (i.e. $s_{(a)b}$ and $s_{(a)d}$, respectively) for the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ that describe the 4D BRST-quantized combined field-theoretic system of the free Abelian 3-form and 1-form gauge theories. In addition, we have given a great deal of importance to the existence of a couple of discrete duality symmetry transformations (in the context of our 4D BRST-quantized theory) and have highlighted their roles in the algebraic structures that are obeyed by the four off-shell nilpotent, infinitesimal and continuous symmetry transformation operators that exist for the coupled (but equivalent) Lagrangian densities of our theory. To be precise, we have provided the physical realizations [cf. Eqs. (14),(17)] of the well-known mathematical relationship $\delta=\pm\,*\,d\,*$ between the nilpotent (i.e. $d^{2}=0,\delta^{2}=0$) (co-)exterior derivatives $(\delta)d$ of differential geometry in terms of the interplay between (i) the discrete duality symmetry transformation operators [cf. Eq. (13)], and (ii) the infinitesimal, continuous and off-shell nilpotent symmetry transformation operators [cf. Eqs. (3),(4),(9),(10)].
@@
Out of the total six anticommutators (that can be constructed from the four nilpotent symmetry operators), we have shown that (i) the two specific anticommutators (i.e. $\{s_{d},\;s_{ab}\}=0$ and $\{s_{ad},\;s_{b}\}=0$) are trivially zero on their own for all the quantum fields of our 4D BRST-quantized theory [except for the fields $\beta_{\mu},\;\bar{\beta}_{\mu},\;\phi_{\mu}$ and $\widetilde{\phi}_{\mu}$ that are present in equation (21) which transform up to a set of U(1) vector gauge symmetry-type transformations], (ii) another two anticommutators (i.e. $\{s_{d},\;s_{ad}\}=0$ and $\{s_{ab},\;s_{b}\}=0$) are zero provided we use specific set of three (out of the total four) CF-type restrictions on our theory, and (iii) the rest of the two anticommutators (i.e. $\{s_{d},\;s_{b}\}=s_{\omega}$ and $\{s_{ab},\;s_{ad}\}=s_{\bar{\omega}}$) define a set of two bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$. We have established, in our present investigation, that there exists a unique bosonic symmetry operator (i.e. $s_{\omega}=-\,s_{\bar{\omega}}$) on the subspace (of the total field space) of quantum fields which is precisely defined by all the four CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) that exist on our 4D BRST-quantized field-theoretic system [cf. Eqs. (27),(31),(32)].
@@
We find that (i) the algebraic structures of the symmetry transformation operators (in our section three and four), and (ii) the ghost number considerations (in the context of off-shell nilpotent symmetry operators) are such that we obtain a two-to-one mapping [i.e. $(s_{b},\;s_{ad})\Rightarrow d,\;(s_{d},\;s_{ab})\Rightarrow\delta,\;(s_{\omega},\;s_{\bar{\omega}})\Rightarrow\Delta$] between (i) the symmetry transformation operators, and (ii) the de Rham cohomological operators of differential geometry (where the discrete duality symmetry operators [cf. Eq. (13)] provide the physical realization(s) of the Hodge duality * operator). The study of the BRST-quantized field-theoretic systems has been mathematically and physically useful in the sense that we have been able to establish that the 2D (non-)Abelian gauge theories (without any interaction with the matter fields) provide [17] a new type of topological field theory (TFT) that captures a few aspects of the Witten-type TFT [18] and some salient features of the Schwarz-type TFT [19]. In addition, such studies have led to the existence of “exotic” fields with the negative kinetic terms (see, e.g. [15] and references therein) that are a set of possible candidates for (i) the “phantom” and/or “ghost” fields of the cyclic, bouncing and self-accelerated cosmological models of the Universe (see, e.g. [20-22]), and (ii) the dark energy/dark matter (see, e.g. [23,24]).
@@
In our future endeavor, we plan to derive (i) all the Noether conserved charges corresponding to the continuous symmetry transformations of our 4D BRST-quantized theory, and (ii) the algebraic structures satisfied by the appropriate conserved charges and study the mapping between the appropriate conserved charges and the de Rham cosmological operators of differential geometry. In other words, we wish to prove that (i) our present 4D BRST-quantized field-theoretic system, and (ii) the Stückelberg-modified massive version of our present system, are a set of the tractable examples for Hodge theory.

### v2

@@ 4 Bosonic Symmetry Transformations: Uniqueness Property and Full Set of CF-Type Restrictions
A close look at the above equation demonstrates that all the individual terms on the r.h.s. incorporate into themselves one of the CF-type restrictions (out of the total four) that exists on our theory. This statement is true for (i) the terms that are within the square bracket of the total spacetime derivative, and (ii) the terms that exist outside the total spacetime derivative. Thus, ultimately, we have been able to establish the uniqueness of the bosonic symmetry transformation operator at the level of symmetry transformations on (i) the individual fields of our theory [cf. Eq. (27)], and (ii) the coupled (but equivalent) Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ [cf. Eqs. (33),(34)]. We would like to repeat, once again, that this key observation is true only on the subspace of quantum fields where all the four CF-type restrictions are satisfied together [cf. Eqs. (27),(33),(34) for more details].
It is interesting to end this section with the final remark that the unique (i.e. $s_{\omega}=-\,s_{\bar{\omega}}$) bosonic symmetry transformation operator commutes (i.e. $s_{\omega},\;s_{(a)b}]=0,\;[s_{\omega},\;s_{(a)d}]=0$) with all the four nilpotent symmetry transformation operators (i.e. $s_{(a)b},\;s_{(a)d}$) of our theory. Thus, ultimately, we have been able to derive the following algebraic structures in the mathematical language of symmetry transformation operators, namely;
$s_{(a)b}^{2}=0,\qquad\;\;s_{(a)d}^{2}=0,\qquad\;\;s_{(a)d}=\pm\;*\;s_{(a)b}\;*,\qquad\;\;s_{\omega}+s_{\bar{\omega}}=0,$
$\{s_{b},\;s_{ad}\}=0,\qquad\{s_{d},\;s_{ab}\}=0,\qquad\{s_{b},\;s_{ab}\}=0,\quad\{s_{d},\;s_{ad}\}=0,$
$\{s_{b},\;s_{d}\}=s_{\omega},\quad\{s_{ab},\;s_{ad}\}=s_{\bar{\omega}},\quad[s_{\omega},\;s_{(a)b}]=0,\quad[s_{\omega},\;s_{(a)d}]=0.$ (35)
A careful look at the above equation establishes that we have obtained the above algebraic structures that are reminiscent of the Hodge algebra (cf. second footnote in Sec. 1) that are satisfied by the de Rham cohomological operators (cf. Sec. 5 for more discussions). Some of the key observations, at this juncture, are (i) the operator relationship: $s_{\omega}+s_{\bar{\omega}}=0$ is valid if and only if all the four CF-type restrictions are invoked together, (ii) the absolute anticommutativity properties: $\{s_{b},\;s_{ab}\}=0$ and $\{s_{d},\;s_{ad}\}=0$ are true only when the specific set of three (out of the total four) CF-type restrictions are imposed from outside [cf. Eqs. (18),(19)], and (iii) the anticommutators $\{s_{b},\;s_{ad}\}=0$ and $\{s_{b},\;s_{ab}\}=0$ are true only up to the U(1) vector gauge symmetry-type transformations [cf. Eq. (21) for details].
5 Summary and Future Perspective
@@
In our present endeavor, we have very briefly sketched the existence of the infinitesimal, continuous and off-shell nilpotent (i.e. $s_{(a)b}^{2}=0,\;s_{(a)d}^{2}=0$) versions of the (anti-)BRST and (anti-)co-BRST symmetry transformation operators (i.e. $s_{(a)b}$ and $s_{(a)d}$, respectively) for the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ that describe the 4D BRST-quantized combined field-theoretic system of the free Abelian 3-form and 1-form gauge theories. We would like to lay emphasis on the fact that this combined field-theoretic system is essential for the existence of the nilpotent (anti-)co-BRST symmetry transformations. We have given a great deal of importance to the existence of a couple of discrete duality symmetry transformations (in the context of our 4D BRST-quantized theory) and have highlighted their roles in the algebraic structures that are obeyed by the four off-shell nilpotent, infinitesimal and continuous symmetry transformation operators and the inter-relationships that exist amongst them [cf. Eqs. (14),(17),(35)]. To be precise, we have provided the physical realizations [cf. Eqs. (14),(17)] of the well-known mathematical relationship $\delta=\pm\,*\,d\,*$ between the nilpotent (i.e. $d^{2}=0,\delta^{2}=0$) (co-)exterior derivatives $(\delta)d$ of differential geometry in terms of the interplay between (i) the discrete duality symmetry transformation operators [cf. Eq. (13)], and (ii) the infinitesimal, continuous and off-shell nilpotent symmetry transformation operators [cf. Eqs. (3),(4),(9),(10)].
Against the backdrop of the above paragraph, we would like to emphasize that the relationships in equations (14) and (17) corroborate our observations that the nilpotent (i.e. $s_{(a)d}^{2}=0,\;s_{(a)b}^{2}=0$) (anti-)co-BRST symmetry transformations $s_{(a)d}$ [cf. Eqs. (9),(3)] and (anti-)BRST symmetry transformations $s_{(a)b}$ [cf. Eqs. (10),(4)] are dual to one-another and the precise choices of the ($\pm$) signs in their relationships [cf. Eqs. (14),(17)] are dictated by the set rules [cf. Eq. (15)] of duality-invariant Abelian $p$-form ($p=1,2,3$...) gauge theories that has been propounded in [19]. One of the highlights of our present investigation is the observation that the reciprocal relationships: $s_{(a)b}=\pm\;*\;s_{(a)d}\;*$ also exist in our 4D BRST-quantized theory. However, the choices of the ($\pm$) signs are not dictated by any kinds of set rules. For the sanctity and the proof of these relationships (i.e. $s_{(a)b}=\pm\;*\;s_{(a)d}\;*$), one has to make the choices for the ($\pm$) signs judiciously. A close look at the nilpotent transformations [cf. Eqs. (3),(4),(9),(10)] demonstrate that the pair of nilpotent operators ($s_{b},\;s_{ad}$) raise the ghost number of a field by one. On the other hand, the nilpotent pair of operators ($s_{d},\;s_{ab}$) lower the ghost number of a field by one. These observations are just like the operations of the exterior and co-exterior derivatives of differential geometry on a given form. Thus, we find that (i) the algebraic structures of the symmetry transformation operators (cf. Secs. 3-5), and (ii) the ghost number considerations (in the context of the off-shell nilpotent symmetry operators) are such that we obtain a two-to-one mapping [i.e. $(s_{b},\;s_{ad})\Rightarrow d,\;(s_{d},\;s_{ab})\Rightarrow\delta,\;(s_{\omega},\;s_{\bar{\omega}})\Rightarrow\Delta$] between (i) the symmetry transformation operators, and (ii) the de Rham cohomological operators (where the discrete duality symmetry operators [cf. Eq. (13)] provide the physical realization(s) of the Hodge duality * operator in the mathematical relationship: $\delta=\pm\,*\,d\,*$ between the nilpotent (i.e. $\delta^{2}=0,\;d^{2}=0$) (co-)exterior derivatives $(\delta)d$).
@@
Out of the total six anticommutators (that can be constructed from the four nilpotent symmetry operators), we have shown that (i) the two specific anticommutators (i.e. $\{s_{d},\;s_{ab}\}=0$ and $\{s_{ad},\;s_{b}\}=0$) are trivially zero on their own for all the quantum fields of our 4D BRST-quantized theory except for the fields $\beta_{\mu},\;\bar{\beta}_{\mu},\;\phi_{\mu}$ and $\widetilde{\phi}_{\mu}$ which transform up to the U(1) vector gauge symmetry-type transformations [cf. Eq. (21)], (ii) another two anticommutators (i.e. $\{s_{d},\;s_{ad}\}=0$ and $\{s_{ab},\;s_{b}\}=0$) are zero provided we use the specific set of three (out of the total four) CF-type restrictions on our theory, and (iii) the rest of the two anticommutators (i.e. $\{s_{d},\;s_{b}\}=s_{\omega}$ and $\{s_{ab},\;s_{ad}\}=s_{\bar{\omega}}$) define a set of two bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$. We have established, in our present investigation, that there exists a unique bosonic symmetry operator (i.e. $s_{\omega}=-\,s_{\bar{\omega}}$) on the subspace of quantum fields which is precisely defined by all the four CF-type restrictions (i.e. $B_{\mu\nu}+\bar{B}_{\mu\nu}=(\partial_{\mu}\phi_{\nu}-\partial_{\nu}\phi_{\mu}),\;{\cal B}_{\mu\nu}+\bar{\cal B}_{\mu\nu}=(\partial_{\mu}\widetilde{\phi}_{\nu}-\partial_{\nu}\widetilde{\phi}_{\mu}),\;f_{\mu}+F_{\mu}=\partial_{\mu}C_{1},\;\bar{f}_{\mu}+\bar{F}_{\mu}=\partial_{\mu}\bar{C}_{1}$) that exist on our 4D BRST-quantized field-theoretic system [cf. Eqs. (27),(33),(34)]. It is worthwhile to point out that the antiommutators $\{s_{b},\;s_{ad}\}$ and $\{s_{ab},\;s_{d}\}$ do not define (cf. Appendix A) a new set of true bosonic symmetry transformations because they (i) change the ghost numbers of the specific fields $\beta_{\mu},\;\bar{\beta}_{\mu},\;\phi_{\mu}$ and $\widetilde{\phi}_{\mu}$ of our 4D BRST-quantized theory by ± $\pm\,2$, (ii) transform these specific fields only up to the U(1) vector gauge symmetry-type transformations, and (iii) disappear in the limit $B_{2}=B_{3}=B_{4}=B_{5}=0$ (which is not the case with the true bosonic symmetry transformations $s_{\omega}$ and $s_{\bar{\omega}}$ that do not change the ghost number of any arbitrary field of our theory when they operate on it [cf. Eqs. (24),(26)]). Thus, for all practical purposes, these anticommutators (i.e. $\{s_{d},\;s_{ad}\}=0$ and $\{s_{ab},\;s_{b}\}=0$) are zero modulo the U(1) vector gauge symmetry-type transformations.
@@
The study of the BRST-quantized field-theoretic systems, as the tractable examples for Hodge theory, has been mathematically and physically useful in the sense that we have been able to establish that the 2D (non-)Abelian gauge theories (without any interaction with the matter fields) provide a new type of topological field theory (TFT) that captures [20] a few aspects of the Witten-type TFTs [21] and some salient features of the Schwarz-type TFTs [22]. In addition, such studies have led to the existence of “exotic” fields with the negative kinetic terms (see, e.g. [18],[23],[24] and references therein) that are a set of possible candidates for (i) the “phantom” and/or “ghost” fields of the cyclic, bouncing and self-accelerated cosmological models of the Universe (see, e.g. [25-27]and references therein), and (ii) the dark energy/dark matter (see, e.g. [28,29] and references therein).
@@
In our future endeavor, we plan to derive (i) all the Noether conserved charges corresponding to the continuous symmetry transformations of our 4D BRST-quantized theory (that incorporate into themselves the ghost-scale symmetry transformations, too), and (ii) the algebraic structures satisfied by the appropriate conserved charges and study the mapping(s) between the appropriate conserved charges and the de Rham cosmological operators of differential geometry. A very interesting future direction is to study the 4D BRST-quantized field-theoretic system of the Stückelberg-modified massive version of our present theory and discuss its discrete and continuous symmetry transformations (and corresponding conserved charges). In other words, we envisage to prove that (i) our present 4D BRST-quantized field-theoretic system, and (ii) the Stückelberg-modified massive version of our present field-theoretic system, are a set of tractable examples for Hodge theory [30].
Declarations
@@
Conflicts of interest
The author declares that there are no conflicts of interest.
Appendix A: Comments on the Anticommutators (20) and (21)
The central objective of our present Appendix is to prove explicitly that the anticommutators $\{s_{b},\;s_{ad}\}\equiv s_{\omega}^{(1)}$ and $\{s_{d},\;s_{ab}\}\equiv s_{\omega}^{(2)}$ do not define a new set of bosonic symmetry transformations in the true sense of the word because they raise/lower the ghost number by two for a field on which they operate). This observation is in contradiction to the fundamental definition of a set of true bosonic symmetry transformations [cf. Eqs. (24),(26)]. The latter do not change the ghost number of a field on which they operate. To corroborate this statement, first of all, we observe the following
$s_{\omega}^{(1)}\bar{\beta}_{\mu}=\partial_{\mu}\big(B_{2}-B_{3}\big),\qquad s_{\omega}^{(1)}\phi_{\mu}=-\,\partial_{\mu}B_{4},\qquad s_{\omega}^{(1)}\widetilde{\phi}_{\mu}=-\,\partial_{\mu}B_{4},$
$s_{\omega}^{(2)}\beta_{\mu}=\partial_{\mu}\big(B_{2}+B_{3}\big),\qquad s_{\omega}^{(2)}\phi_{\mu}=+\,\partial_{\mu}B_{5},\qquad s_{\omega}^{(2)}\widetilde{\phi}_{\mu}=-\,\partial_{\mu}B_{5}.$ (A.1)
which correspond to the transformations listed in our earlier equation (21). Under the above infinitesimal transformations, we note the following
$s_{\omega}^{(1)}\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\Big[B_{4}\,\partial^{\mu}\big(B_{2}-B_{3}\big)-\big(B_{2}-B_{3}\big)\,\partial^{\mu}B_{4}\Big]\equiv s_{\omega}^{(1)}\,{\cal L}_{(\bar{B})},$
$s_{\omega}^{(2)}\,{\cal L}_{(B)}=\dfrac{1}{2}\,\partial_{\mu}\Big[\big(B_{2}+B_{3}\big)\,\partial^{\mu}B_{5}-B_{5}\,\partial^{\mu}\big(B_{2}+B_{3}\big)\Big]\equiv s_{\omega}^{(2)}\,{\cal L}_{(\bar{B})},$ (A.2)
which demonstrate that the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ transform to the total spacetime derivatives thereby rendering the action integral $S=\int d^{4}x\,{\cal L}_{(B)}\equiv\int d^{4}x\,{\cal L}_{(\bar{B})}$ invariant (i.e. $s_{\omega}^{(1)}S=0,\;s_{\omega}^{(2)}S=0$) due to Gauss’s divergence theorem (because of which all the physical fields vanish off as $x\to\pm\,\infty$). Thus, we conclude, ultimately, that the continuous and infinitesimal transformations, [cf. Eq. (A.1)], corresponding to the bosonic operators $s_{\omega}^{(1)}$ and $s_{\omega}^{(2)}$, are the symmetry transformations of the action integral.
We end this short (but very important) Appendix with the following crucial remarks. First of all, we note that the true bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ [cf. Eqs. (24),(26)] do not change the ghost number of the individual fields of the coupled Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ when they operate on them. Second, the (anti-)ghost fields of the Lagrangian densities ${\cal L}_{(B)}$ and ${\cal L}_{(\bar{B})}$ either do not transform at all or transform up to the U(1) gauge symmetry-type transformations under $s_{\omega}$ and $s_{\bar{\omega}}$ [cf. Eqs. (24),(26)]. Third, we observe that the symmetry transformation operator $s_{\omega}^{(1)}$ raises the ghost number of a field by two when it operates on it [cf. Eq. (A.1)] which implies that this transformation operator is not like the true bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ [cf. Eqs. (24),(26)]. Fourth, it is interesting to point out that the symmetry operator $s_{\omega}^{(2)}$ lowers the ghost number by two when it operates on a field [cf. Eq. (A.1)]. Hence, this operator is also not like the true bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ [cf. Eqs. (24),(26)]. Fifth, under the symmetry transformation operators $s_{\omega}^{(1)}$ and $s_{\omega}^{(2)}$ only the bosonic vector (anti-)ghost fields $(\bar{\beta}_{\mu})\beta_{\mu}$ and the (axial-)vector fields $(\widetilde{\phi}_{\mu})\phi_{\mu}$ transform unlike the transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ [cf. Eqs. (24),(26)] under which the gauge fields as well as the fermionic and bosonic (anti-)ghost fields transform. Sixth, it is straightforward to note that in the limit: $B_{2}=B_{3}=B_{4}=B_{5}=0$, the symmetry transformation operators $s_{\omega}^{(1)}$ and $s_{\omega}^{(2)}$ disappear completely. However, the the true bosonic symmetry transformation operators $s_{\omega}$ and $s_{\bar{\omega}}$ [cf. Eqs. (24),(26)] survive in this limits, too, along with the four off-shell nilpotent symmetry transformation operators [cf. Eqs. (3),(4),(9),(10)]. Finally, it is clear that if the nilpotent symmetry transformation operators are identified with the nilpotent (co-)exterior derivatives, the Laplacian operator will be identified with (i) the unique true bosonic symmetry transformation operator (e.g. $s_{\omega}$) which does not change the ghost number of the field on which it operates, and (ii) the other (ghost number changing) bosonic symmetry transformation operators (which have been listed in (A.1)) can not be, obviously, identified with the Laplacian operator.